# Soda-Eval: Open-Domain Dialogue Evaluation in the age of LLMs

### v1

@@ 1 Introduction
Current evaluation practices of dialogue systems are still highly dependent on human evaluation. Whilst evaluation using strong LLMs has increased in popularity, human evaluation remains the gold standard since it offers a direct view towards human preferences and satisfaction in the interaction, which is the major goal of open domain dialogue. Leaderboard platforms such as ChatBot Arena (Chiang et al., 2024) rank models by calculating an Elo rating obtained from pairwise comparisons of chatbot responses. However, direct assessments provide a more granular evaluation of response quality that pairwise comparisons lack, especially when comparing models that differ only slightly in quality but are otherwise similar (Smith et al., 2022).
Despite the potential benefits of direct assessments for open-domain dialogues, the evaluation community is constrained to using a limited number of benchmark datasets, many of which have become outdated. For instance, FED Mehri and Eskenazi (2020a), a typically used benchmark in dialogue evaluation, annotates responses generated by arguably obsolete chatbots, and targets quality aspects such as fluency or relevance. While the annotation of these quality aspects may have been of interest at the time, it is not clear if contemporary chatbots still suffer from these issues.
@@ 2 Related Work
With the widespread introduction of LLMs, this limitation has been mostly circumvented. Most recent contributions typically leverage GPT-3.5-Turbo or GPT-4 for dialogue quality assessments (Liu et al., 2023; Lin and Chen, 2023; Mendonça et al., 2023). Of note, Finch et al. (2023b) investigated the ability of GPT-3.5-Turbo to perform evaluation of real human-bot dialogues using the ABC evaluation framework (Finch et al., 2023a). There have also been some efforts to divest from closed-source LLMs, with XDial-Eval (Zhang et al., 2023) probing the evaluation capabilities of several open source LLMs against GPT-3.5-Turbo (Ouyang et al., 2022). However, to the best of our knowledge, our work is the first that attempts to directly distil evaluation knowledge from closed-source LLMs to smaller open-source ones.
@@ 4.2.1 Issue Detection
We sampled 200 examples from the test set (each annotator validated 100 examples, and we recruited 3 annotators per example) for validation. To reduce annotator confusion, we ensured that the examples that contain issues belong to the first flagged responses of the dialogue. This is because we found human annotators had difficulties annotating a response when earlier responses had issues.
@@ 5.3 In-domain Results
As expected, finetuning yields higher BLEU scores when compared to few-shot prompting, with the weakest model achieving much better scores than all of the instruction tuned models. However, a high BLEU score does not ensure the validity of the explanation. Consequently, we present explanation validation results in Table 7. If we exclude from analysis Engagement as an issue we see that all models produce valid explanations over 60% of the times, even when excluding issue-free responses from consideration. Additionally, we observe that the finetuned models produce more valid explanations than their corresponding instruction-tuned models, across all scenarios. For instance, finetuning Phi-3-mini yields a 15% absolute improvement in the detection of major issues (and 22% for all issues).
@@ 7 Limitations
We acknowledge that the selection of Soda as our base dialogue dataset for analysis and annotation comes with some downsides. Firstly, since the whole interaction is synthetically generated by a single model, it may not accurately represent dyadic conversational dynamics. For instance, we expect a human would point out a major issue in the response directly, instead of continuing the conversation unimpeded (Petrak et al., 2023). However, since we conduct a turn level assessment, this issue is mostly circumvented, at least for the initial turn the issue is presented.
@@ Appendix A Dialogue analysis
(Finch et al., 2023a) conducted a comprehensive evaluation of four open-domain chatbots using the ABC-Eval 16 behavior labels. These chatbots are as follows: (1) BlenderBot-DECODE (Nie et al., 2021); (2) BlenderBot2  [^5] BlenderBot 2 model card; (3) BART-FiD-RAG (Shuster et al., 2021); and (4) Emora (Finch et al., 2020).
@@ D.2 Overall Assessment Annotations
Quality control in crowdsourcing is the subject of significant research in the literature (Daniel et al., 2018). While human evaluators bring subjective insight in to the assessment, they may also be prone to overlooking subtle problems or inconsistencies within the conversation, especially when they are under significant cognitive load. As such, we did an experiment that attempts to understand if integrating LLMs as an assistant in the evaluation process can help annotators accurately rate dialogue responses.
@@ Appendix E Implementation Details
We train our models with a language generation objective. Provided with a dialogue history $c$ and a candidate response $r$, they are tasked to output, in natural language, an overall assessment of the response and a corresponding score $s\in[1,5]$. We finetune our models on a single RTX A6000 48GB GPU or A100 80GB GPU using Huggingface Transfomers with TRL Supervised Fine-tuning Trainer (SFT) 777huggingface.co/docs/trl/sft_trainer and PEFT888huggingface.co/docs/peft for 3 epochs. We conduct a single finetuning run from the base instruction models (full precision) using LoRA (Hu et al., 2021), with $r=8$, $\alpha=32$ and dropout set to 0.1. Gradient accumulation steps is set to 4 with a learning rate of $41e-$ 4. Batch size was set to maximise VRAM consumption, ranging from 4 up to 64 per device.

### v3

@@ 1 Introduction
Current evaluation practices of dialogue systems are still highly dependent on human evaluation. Whilst evaluation using strong LLMs has increased in popularity, human evaluation remains the gold standard since it offers a direct view towards human preferences and satisfaction in the interaction, which is the major goal of open domain dialogue. Leaderboard platforms such as ChatBot Arena (Chiang et al., 2024) rank models by calculating an Elo rating obtained from pairwise comparisons of chatbot responses. However, DA (direct assessments) of responses provide a more granular evaluation of response quality that pairwise comparisons lack, especially when comparing models that differ only slightly in quality but are otherwise similar (Smith et al., 2022).
Despite the potential benefits of direct assessments for open-domain dialogues, the evaluation community is constrained to using a limited number of benchmark datasets, many of which have become outdated (Mendonça et al., 2024). For instance, FED Mehri and Eskenazi (2020a), a typically used benchmark in dialogue evaluation, annotates responses generated by arguably obsolete chatbots, and targets quality aspects such as fluency or relevance. While the annotation of these quality aspects may have been of interest at the time, it is not clear if contemporary chatbots still suffer from these issues.
@@ 2 Related Work
With the widespread introduction of LLMs, this limitation has been mostly circumvented. Most recent contributions typically leverage GPT-3.5-Turbo or GPT-4 for dialogue quality assessments (Liu et al., 2023; Lin and Chen, 2023; Mendonça et al., 2023). Of note, Finch et al. (2023b) investigated the ability of GPT-3.5-Turbo to perform evaluation of real human-bot dialogues using the ABC evaluation framework (Finch et al., 2023a). There have also been some efforts to divest from closed-source LLMs, with XDial-Eval (Zhang et al., 2023) probing the evaluation capabilities of several open source LLMs against GPT-3.5-Turbo (Ouyang et al., 2022) and Mendonca et al. (2024) proposing an explainable evaluator of coherence by synthetically generating positive and adversarial negative responses using a closed-source LLM to finetune a smaller, open-source one. However, to the best of our knowledge, our work is the first that conducts direct knowledge distillation of multiple dialogue quality aspects.
@@ 4.2.1 Issue Detection
We sampled 200 examples [^5] This validation sample size is in line with other works (Zeng et al., 2024; Wang et al., 2024). from the test set (each annotator validated 100 examples, and we recruited 3 annotators per example) for validation. To reduce annotator confusion, we ensured that the examples that contain issues belong to the first flagged responses of the dialogue. This is because we found human annotators had difficulties annotating a response when earlier responses had issues.
@@ 5.3 In-domain Results
As expected, finetuning yields higher BLEU scores when compared to few-shot prompting, with the weakest model achieving much better scores than all of the instruction tuned models. However, a high BLEU score does not ensure the validity of the explanation. Consequently, we present explanation validation results in Table 7. If we exclude from analysis Engagement as an issue we see that all models produce valid explanations over 60% of the times, even when excluding issue-free responses from consideration. Additionally, we observe that the finetuned models produce more valid explanations than their corresponding instruction-tuned models, across all scenarios. For instance, finetuning Phi-3-mini yields a 15% absolute improvement in the detection of major issues.
When including Engagement in the analysis, we note that the models struggle with identifying engagement issues (as reported by GPT-4), with all open-access models achieving under 30% validity. The exception to this trend is GPT-3.5-Turbo – this can be explained by the fact it belongs to the same family as GPT-4, and as such is likely to be trained using similar data. When finetuning the models, we observe an even higher improvement when compared to the response set without engagement issues (e.g., we report an absolute improvement of 22% validity with Phi-3-mini), which indicates the finetuning step has led the models to better align themselves with the subjective assessments of the teacher model.
@@ 7 Limitations
We acknowledge that the selection of Soda as our base dialogue dataset for analysis and annotation comes with some downsides. Firstly, given the dataset is generated from a single source (social commonsense distilled with GPT-3.5), its dialogues may not fully capture the diversity of human-chatbot interactions. Secondly, since the whole interaction is synthetically generated in a single forward pass, it may not accurately represent dyadic conversational dynamics. For instance, we expect a human would point out a major issue in the response directly, instead of continuing the conversation unimpeded (Petrak et al., 2023). However, since we conduct a turn level assessment, this issue is mostly circumvented, at least for the initial turn the issue is presented.
@@ Appendix A Dialogue analysis
Finch et al. (2023a) conducted a comprehensive evaluation of four open-domain chatbots using the ABC-Eval 16 behavior labels. These chatbots are as follows: (1) BlenderBot-DECODE (Nie et al., 2021); (2) BlenderBot2  [^6] BlenderBot 2 model card; (3) BART-FiD-RAG (Shuster et al., 2021); and (4) Emora (Finch et al., 2020).
@@ D.2 Overall Assessment Annotations
Quality control in crowdsourcing is the subject of significant research in the literature (Daniel et al., 2018). While human evaluators bring subjective insight in to the assessment, they may also be prone to overlooking subtle problems or inconsistencies within the conversation, especially when they are under significant cognitive load. As such, we conducted an experiment that attempts to understand if integrating LLMs as an assistant in the evaluation process can help annotators accurately rate dialogue responses.
@@ Appendix E Implementation Details
We train our models with a language generation objective. Provided with a dialogue history $c$ and a candidate response $r$, they are tasked to output, in natural language, an overall assessment of the response and a corresponding score $s\in[1,5]$. We finetune our models on a single RTX A6000 48GB GPU or A100 80GB GPU using Huggingface Transfomers with TRL Supervised Fine-tuning Trainer (SFT) 888huggingface.co/docs/trl/sft_trainer and PEFT999huggingface.co/docs/peft for 3 epochs. We conduct a single finetuning run from the base instruction models (full precision) using LoRA (Hu et al., 2021), with $r=8$, $\alpha=32$ and dropout set to 0.1. Gradient accumulation steps is set to 4 with a learning rate of $41e-$ 4. Batch size was set to maximise VRAM consumption, ranging from 4 up to 64 per device.